A soldering iron has a cylindrical tip of 2.5 mm in diameter and 20 mm in length. With 

age and usage, the tip has oxidized and has an emissivity of [a]. Assuming that the 

average convection heat transfer coefficient over the soldering iron tip is [b] W/m2

·K, 

and the surrounding air temperature is 20°C, determine the power required to 

maintain the tip at 400°C. 


1
Expert's answer
2021-05-19T06:07:31-0400

Assume that the emissivity is a=0.8 and the convection heat transfer coefficient is h=25 W/m2.

Find the surface area of the cylindrical tip:


"A=\pi DL+\pi\frac{D^2}{4}=\pi D\bigg(L+\frac D4\bigg)."

The heat transfer through convection (in the following equation "T_r, T_s" are required and surrounding temperatures respectively):


"Q_c=hA(T_r-T_s),\\
Q_c=h\pi D\bigg(L+\frac D4\bigg)(T_r-T_s),\\\space\\
Q_c=1.53\text{ W}."

The heat transfer through radiation can be found by Stephan-Boltzman law:


"Q_r=\epsilon\sigma A(T_r^4-T_s^4),\\
Q_r=\epsilon\sigma\pi D\bigg(L+\frac D4\bigg)(T_r^4-T_s^4)=1.45\text{ W}."

The total heat power required:


"Q=Q_c+Q_r=2.98\text{ W}."